A4. Using the Chebyshev's inequality, nd the value of k that will guarantee that the
probability is 0:95 when the deviation of X from its mean is not more than k.
1
Expert's answer
2020-06-08T20:26:54-0400

Chebyshev's inequality


"P(|X-\mu|\geq k\sigma)\leq{1\over k^2}"

Thus, the probability that a random variable takes a value more than k standard deviations away from its mean is at most "1\/k^2."


"k=4: {1\over k^2}={1\over 16}=0.0625(6.25\ \%)"

"100\ \%-6.25 \ \%=93.75\ \%<95\ \%"

"k=5: {1\over k^2}={1\over 25}=0.04(4\ \%)"

"100\ \%-4 \ \%=96\ \%>95\ \%"

"k=5."